What is the solution of |2x - 1| = 5 ?
2x-1=5
2x=5+1
2x=6
2/2=6/2
X=3
|2x-1|=-5
2x=5+1
2x=-4
2/2=-4/2
x=-2
so (x=3 and -2)

Found 2 solutions by philline_palana, cahilldeb:
Answer by philline_palana(20)

We have two answers for your problem because it is an absolute value,
|2x - 1| = 5
first solution:
|2x - 1| = 5 >>remove the absolute value sign
2x - 1 = 5 >>transpose
2x = 5 + 1 >>add
2x = 6 >>divide the equation by two
2x/2 = 6/2
x = 3 >>we are not yet done
second solution:
|2x - 1| = 5 >>remove the absolute value sign and change the sign of 5 into negative
2x - 1 = -5 >>transpose
2x = -5 + 1 >>add
2x = -4 >>divide the equation by two
2x/2 = -4/2
x = -2 >>we are done.
so for this problem, our two solutions are: x=3 and x=-2

Answer by cahilldeb(1)

|2x-1|=5
2x=5+1
2x=6
2x/2=6/2
x=3
|2x-1|=5
2x-1=-5
2x=-4
2x/2=-4/2
x=-2
answer is (x=3 and -2)